OpenStudy (anonymous):

3 sin x=2 cos^2 x solve

OpenStudy (anonymous):

try like this: since cos^2x=1-sin^2x 3sinx=2(1-sin^2x) or 2sin^2x+3sinx-2=0 now let sinx=t, then 2t^2+3t-2=0 and solve for t. Later you will have to find corresponding values of x

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw